# Y. Grandati1 and C. Quesne2

### v1

@@ A One-step DT
Any formal eigenfunction $\psi_{\nu}(x)$ of $V(x)$(i.e., of $\widehat{H}$) can be used as a seed function for a Darboux transformation (DT) $A(\psi_{\nu})$, which associates with the potential $V(x)$ a modified potential
@@ A Spectrum
The trigonometric Darboux-Pöschl-Teller (TDPT) potential (with zero ground-state energy) is defined on $]0,\pi/2[$ by
@@ IV ONE-STEP REGULAR EXTENSIONS OF THE TDPT POTENTIAL
For $\max(N,M)\leq n<N+M$, we have $E_{-n-1}(N,M)<0$ and $\psi_{-n-1}\left(z;N,M;\lambda\right)$ is disconjugated on $x\in$ $]0,\pi/2[$ $(z\in$ ]1,+1[)]-1,+1[)] - 1 , + 1 [ ) and consequently admits at most one zero on this interval. As proven in [24], $\psi_{-n-1}$ has no node there iff the parameters satisfy the following regularity conditions:
@@
With the choices above, we can then use $\psi_{-n-1}$ as seed function of a one-step DT to produce a rational extension of the TDPT potential, which is perfectly regular on $]0,\pi/2[$. This extended potential is given by [24]
@@
Its spectrum is given by $E_{k}\left(N,M\right)$, $k\in\left\{-n-1,0,1,...,\right\}$, with the corresponding eigenstates:
@@
×$\times\left|\begin{array}[]{cc}p_{n}^{(-N,-M)}(z;\lambda)&(1-z)^{N}(1+z)^{M}P_{k}^{(N,M)}(z)\\ np_{n-1}^{\left(-N+1,-M+1\right)}\left(z;a_{n}^{(N,M)}\lambda\right)&-2(k+1)(1-z)^{N-1}(1+z)^{M-1}P_{k+1}^{(N-1,M-1)}(z)\end{array}\right|$ (4.6)
@@
The set of $Q_{k}^{(n)}(z;N,M;\lambda)$, $k=-n-1,0,1,2,\ldots$, is made of orthogonal polynomials on ]1,1[]-1,1[] - 1 , 1 [ with respect to the measure
@@ A Regularity conditions
We are interested in chains for which both extensions $V^{\left(-n_{1}-1\right)}$ and $V^{\left(-n_{1}-1,-n_{2}-1\right)}$ are regular. For such a purpose, we have first to choose $\max\left(N,M\right)\leq n_{1}<N+M$ and $\lambda_{1}$ in order to satisfy the regularity condition (4.1), which ensures that the seed function $\psi_{-n_{1}-1}$ has no node on $]0,\pi/2[$. We then choose $n_{2}$ in order that the seed function of the second DT $\psi_{-n_{2}-1}^{\left(-n_{1}-1\right)}$ be in the disconjugacy sector of $V^{\left(-n_{1}-1\right)}$, namely
@@
$V^{\left(-n_{1}-1,-n_{2}-1\right)}\left(x;N,M;\lambda_{1},\lambda_{2}\right)$ is then regular when $\psi_{-n_{2}-1}^{\left(-n_{1}-1\right)}\left(x;N,M;\lambda_{1},\lambda_{2}\right)$ is nodeless on $]0,\pi/2[$. Since $\psi_{-n_{1}-1}$ does not change its sign on this interval, this corresponds to
@@
×$\times W\left(p_{n_{1}}^{\left(-N,-M\right)}\left(z;\lambda_{1}\right),p_{n_{2}}^{\left(-N,-M\right)}\left(z;\lambda_{2}\right)\mid z\right),$
@@
×$\times\left(E_{-n_{1}-1}\left(N,M\right)-E_{-n_{2}-1}\left(N,M\right)\right).$
@@
×$\times\left(E_{-n_{1}-1}\left(N,M\right)-E_{-n_{2}-1}\left(N,M\right)\right).$
@@
$\left\{\begin{array}[]{c}(M,n_{1}-N,n_{2}-N)\in\left(2\mathbb{N}\right)^{3}:0<\lambda_{1}<\lambda_{n_{1}}^{\left(N,M\right)},\lambda_{2}<0\text{ or }\lambda_{2}>\lambda_{n_{2}}^{(N,M)};\\[5.69046pt] (M,n_{1}-N,n_{2}-N)\in\left(2\mathbb{N}\right)^{2}\times(2\mathbb{N}+1):0<\lambda_{1}<\lambda_{n_{1}}^{\left(N,M\right)},-\lambda_{n_{2}}^{(N,M)}<\lambda_{2}<0;\\[5.69046pt] (M,n_{1}-N,n_{2}-N)\in 2\mathbb{N}\times(2\mathbb{N}+1)\times 2\mathbb{N}:\lambda_{1}<-\lambda_{n_{1}}^{(N,M)}\text{ or }\lambda_{1}>0,\lambda_{2}<0\\ \text{ or }\lambda_{2}>\lambda_{n_{2}}^{(N,M)};\\[5.69046pt] (M,n_{1}-N,n_{2}-N)\in 2\mathbb{N}\times(2\mathbb{N}+1)^{2}:\lambda_{1}<-\lambda_{n_{1}}^{(N,M)}\text{ or }\lambda_{1}>0,\\ -\lambda_{n_{2}}^{(N,M)}<\lambda_{2}<0;\\[5.69046pt] (M,n_{1}-N,n_{2}-N)\in(2\mathbb{N}+1)\times(2\mathbb{N})^{2}:\lambda_{1}<0\text{ or }\lambda_{1}>\lambda_{n_{1}}^{(N,M)},\\ 0<\lambda_{2}<\lambda_{n_{2}}^{(N,M)};\\[5.69046pt] (M,n_{1}-N,n_{2}-N)\in(2\mathbb{N}+1)\times 2\mathbb{N}\times(2\mathbb{N}+1):\lambda_{1}<0\text{ or }\lambda_{1}>\lambda_{n_{1}}^{(N,M)},\\ \lambda_{2}<-\lambda_{n_{2}}^{(N,M)}\text{ or }\lambda_{2}>0;\\[5.69046pt] (M,n_{1}-N,n_{2}-N)\in(2\mathbb{N}+1)^{2}\times 2\mathbb{N}:-\lambda_{n_{1}}^{(N,M)}<\lambda_{1}<0,0<\lambda_{2}<\lambda_{n_{2}}^{(N,M)};\\[5.69046pt] (M,n_{1}-N,n_{2}-N)\in(2\mathbb{N}+1)^{3}:-\lambda_{n_{1}}^{(N,M)}<\lambda_{1}<0,\lambda_{2}<-\lambda_{n_{2}}^{(N,M)}\text{ or }\lambda_{2}>0.\end{array}\right.$ (5.27)
@@ B Rationally-extended potential
For $V^{(-n_{1}-1,-n_{2}-1)}(x;N,M;\lambda_{1},\lambda_{2})$ to be a confining potential on $]0,\pi/2[$, we need to impose this property to $V(x;N-2,M-2)$. This is achieved for $N,M\geq 3$. For such a choice and provided $n_{1}$ and $n_{2}$ satisfy Eq. (5.4), $V^{(-n_{1}-1,-n_{2}-1)}(x;N,M;\lambda_{1},\lambda_{2})$ is strongly repulsive in both 0 and $\pi/2$, so that in each extremity only one basis solution is quadratically integrable and the corresponding Hamiltonian is essentially self-adjoint [33].
@@
×$\qquad\times p_{n_{2}-1}^{(-N+1,-M+1)}\left(z;a_{n_{2}}^{(N,M)}\lambda_{2}\right).$ (5.42)
Due to the orthogonality properties of the $\psi_{k}^{(-n_{1}-1,-n_{2}-1)}(x;N,M;\lambda_{1},\lambda_{2})$, we deduce that the $Q_{k}^{(n_{1},n_{2})}(z;N,M;\lambda_{1},\lambda_{2})$, $k=-n_{1}-1,-n_{2}-1,0,1,2,\ldots$, are orthogonal polynomials on ]1,1[]-1,1[] - 1 , 1 [ with respect to the measure
@@ C Explicit examples
Consider for instance the case where $N=M=3$. The seed function indices $\left(n_{1},n_{2}\right)$ of the two-step chain can be chosen in the set ($63\leq n_{2}<n_{1}<$ 6), so that $(n_{1},n_{2})\in\left\{\left(4,3\right),(5,4),(5,3)\right\}$. As for the possible values of the regularity parameters $\lambda_{n_{i}}^{\left(N,M\right)},\ i=1,2$, we get from Eq. (3.22)
$\lambda_{3}^{\left(3,3\right)}=3,\ \lambda_{4}^{\left(3,3\right)}=2,\ \lambda_{5}^{\left(3,3\right)}=1/6.$ (5.44)
@@
Figure 1: Potential $V^{(-5,-4)}(z;3,3;-1,1)$ in terms of $z$ for $-1<z<1$.
Figure 2: Potential $V^{(-5,-4)}(z;3,3;t,1)$ in terms of $z$ and $t$ for $-1<z<1$ and $-2<t<0$.
Figure 3: Potential $V^{(-5,-4)}(z;3,3;-1,\mu)$ in terms of $z$ and $\mu$ for $-1<z<1$ and $30<\mu<$ 3.
VI $m$-STEP REGULAR EXTENSIONS OF THE TDPT POTENTIAL
@@ A Regularity conditions
On assuming that the extension obtained at the $(m-1)$th step is regular, which imposes that $\max\left(N,M\right)\leq n_{m-1}<\cdots<n_{2}<n_{1}<N+M$ and definite conditions on $\lambda_{1},\lambda_{2},\ldots,\lambda_{m-1}$, we want to impose that the same is true at the $m$th step. For this to occur, $\psi^{(N_{m-1})}_{-n_{m}-1}$ must be in the disconjugacy sector of $V^{(N_{m-1})}$, i.e., $\max\left(N,M\right)\leq n_{m}<n_{m-1}<\cdots<n_{1}<N+M$, and, in addition, it must not change sign on $]0,\pi/2[$, which leads to the condition
@@
×$\qquad\times W\left(p_{n_{1}}^{(-N,-M)}(z;\lambda_{1}),p_{n_{2}}^{(-N,-M)}(z;\lambda_{2}),\ldots,p_{n_{m}}^{(-N,-M)}(z;\lambda_{m})\mid z\right),$ (6.4)
@@
×$\qquad\times\frac{b_{n_{1}}^{\left(N,M\right)}...b_{n_{m}}^{\left(N,M\right)}}{2^{m\left(m-1\right)/2}\left(M-1\right)^{m-1}\left(M-2\right)^{m-2}...(M-m+1)}D(n_{1},...,n_{m}),$ (6.33)
@@
As we have assumed that the ($m-1$)-step chain $N_{m-1}=\left(-n_{1}-1,\ldots,-n_{{m-1}}-1\right)$ is regular at each level, the condition above then implies that
@@ B Rationally-extended potential
which turns out to be a rational extension of $V(x;N-m,M-m)$,
@@
For $V^{\left(N_{m}\right)}\left(x;N,M;\lambda_{1},\ldots,\lambda_{m}\right)$ to be a confining potential on $]0,\pi/2[$, we need to impose that $V\left(x;N-m,M-m\right)$ has this property, which is achieved for $N,M>m+1/2$, thence $N,M\geq m+1$. For such $N,M$ values, we note that $V^{\left(N_{m}\right)}\left(x;N,M;\lambda_{1},\ldots,\lambda_{m}\right)$ is actually strongly repulsive in both 00 and $\pi/2$, since the singularities are there of the type $g/x^{2}$ ($g\geq 3/4$) and $g/\left(\pi/2-x\right)^{2}$ ($g\geq 3/4$), respectively. This means that at each extremity, only one basis solution is quadratically integrable [33].
@@
×$\qquad\times\frac{W(p_{n_{1}}^{(-N,-M)}(z;\lambda_{1}),\ldots,p_{n_{m}}^{(-N,-M)}(z;\lambda_{m}),(1-z)^{N}(1+z)^{M}P_{k}^{(N,M)}(z)\mid z)}{W\left(p_{n_{1}}^{(-N,-M)}(z;\lambda_{1}),p_{n_{2}}^{(-N,-M)}(z;\lambda_{2}),\ldots,p_{n_{m}}^{(-N,-M)}(z;\lambda_{m})\mid z\right)}$
@@
×$\qquad\times\frac{\psi_{-1}(x;N-m,M-m)}{\det\left[\mathbf{R}\left(z;\lambda_{1},...,\lambda_{m}\right)\right]},$ (6.49)
@@
×$\times\frac{W(p_{n_{1}}^{(-N,-M)}(z;\lambda_{1}),\ldots,\check{p}_{n_{i}}^{(-N,-M)}(z;\lambda_{i}),\ldots,p_{n_{m}}^{(-N,-M)}(z;\lambda_{m})\mid z)}{W\left(p_{n_{1}}^{(-N,-M)}(z;\lambda_{1}),\ldots,p_{n_{m}}^{(-N,-M)}(z;\lambda_{m})\mid z\right)},$ (6.53)
@@
where $Q^{(n_{1},n_{2};\ldots,n_{m})}_{-n_{i}-1}(x;N,M;\lambda_{1},\ldots,\lambda_{m})$ is obtained from the determinant of $\mathbf{R}(z;\lambda_{1},\ldots,\lambda_{m})$ by suppressing the $m^{\rm th}$ row and the $i^{\rm th}$ column, i.e., by a minor that we demote by $\det[\mathbf{R}(z;\lambda_{1},\ldots,\lambda_{m})]^{m,i}$,
@@
Due to the orthogonality properties of the $\psi_{k}^{(N_{m})}$, the polynomials $Q_{k}^{(n_{1},n_{2},\ldots,n_{m})}(z;N,M;\lambda_{1},\ldots,\lambda_{m})$, $k=-n_{1}-1$, …, $-n_{m}-1,0,1,2,\ldots$, form a set of orthogonal polynomials on ]1,1[]-1,1[] - 1 , 1 [ with respect of the measure

### v2

@@ A One-step DT
Any formal eigenfunction $\psi_{\nu}(x)$ of V(x)V(x)(i.e., of $\widehat{H}$) can be used as a seed function for a Darboux transformation (DT) $A(\psi_{\nu})$, which associates with the potential $V(x)$ a modified potential
@@ A Spectrum
The trigonometric Darboux-Pöschl-Teller (TDPT) potential (with zero ground-state energy) is defined on $0,\pi/2[$ by
@@ IV ONE-STEP REGULAR EXTENSIONS OF THE TDPT POTENTIAL
For $\max(N,M)\leq n<N+M$, we have $E_{-n-1}(N,M)<0$ and $\psi_{-n-1}\left(z;N,M;\lambda\right)$ is disconjugated on $x\in$ $0,\pi/2[$ $(z\in$ ]1,+1[)]-1,+1[) and consequently admits at most one zero on this interval. As proven in [24], $\psi_{-n-1}$ has no node there iff the parameters satisfy the following regularity conditions:
@@
With the choices above, we can then use $\psi_{-n-1}$ as seed function of a one-step DT to produce a rational extension of the TDPT potential, which is perfectly regular on $0,\pi/2[$. This extended potential is given by [24]
@@
[…]
@@
$\times\left|\begin{array}[]{cc}p_{n}^{(-N,-M)}(z;\lambda)&(1-z)^{N}(1+z)^{M}P_{k}^{(N,M)}(z)\\ np_{n-1}^{\left(-N+1,-M+1\right)}\left(z;a_{n}^{(N,M)}\lambda\right)&-2(k+1)(1-z)^{N-1}(1+z)^{M-1}P_{k+1}^{(N-1,M-1)}(z)\end{array}\right|$ (4.6)
@@
The set of $Q_{k}^{(n)}(z;N,M;\lambda)$, $k=-n-1,0,1,2,\ldots$, is made of orthogonal polynomials on ]1,1[]-1,1[ with respect to the measure
@@ A Regularity conditions
We are interested in chains for which both extensions $V^{\left(-n_{1}-1\right)}$ and $V^{\left(-n_{1}-1,-n_{2}-1\right)}$ are regular. For such a purpose, we have first to choose $\max\left(N,M\right)\leq n_{1}<N+M$ and $\lambda_{1}$ in order to satisfy the regularity condition (4.1), which ensures that the seed function $\psi_{-n_{1}-1}$ has no node on $0,\pi/2[$. We then choose $n_{2}$ in order that the seed function of the second DT $\psi_{-n_{2}-1}^{\left(-n_{1}-1\right)}$ be in the disconjugacy sector of $V^{\left(-n_{1}-1\right)}$, namely
@@
$V^{\left(-n_{1}-1,-n_{2}-1\right)}\left(x;N,M;\lambda_{1},\lambda_{2}\right)$ is then regular when $\psi_{-n_{2}-1}^{\left(-n_{1}-1\right)}\left(x;N,M;\lambda_{1},\lambda_{2}\right)$ is nodeless on $0,\pi/2[$. Since $\psi_{-n_{1}-1}$ does not change its sign on this interval, this corresponds to
@@
$\times W\left(p_{n_{1}}^{\left(-N,-M\right)}\left(z;\lambda_{1}\right),p_{n_{2}}^{\left(-N,-M\right)}\left(z;\lambda_{2}\right)\mid z\right),$
@@
$\times\left(E_{-n_{1}-1}\left(N,M\right)-E_{-n_{2}-1}\left(N,M\right)\right).$
@@
$\times\left(E_{-n_{1}-1}\left(N,M\right)-E_{-n_{2}-1}\left(N,M\right)\right).$
@@
[…]
@@ B Rationally-extended potential
For $V^{(-n_{1}-1,-n_{2}-1)}(x;N,M;\lambda_{1},\lambda_{2})$ to be a confining potential on $0,\pi/2[$, we need to impose this property to V(x;N2,M2)V(x;N-2,M-2). This is achieved for $N,M\geq 3$. For such a choice and provided $n_{1}$ and $n_{2}$ satisfy Eq. (5.4), $V^{(-n_{1}-1,-n_{2}-1)}(x;N,M;\lambda_{1},\lambda_{2})$ is strongly repulsive in both 0 and $\pi/2$, so that in each extremity only one basis solution is quadratically integrable and the corresponding Hamiltonian is essentially self-adjoint [33].
@@
$\qquad\times p_{n_{2}-1}^{(-N+1,-M+1)}\left(z;a_{n_{2}}^{(N,M)}\lambda_{2}\right).$ (5.42)
Due to the orthogonality properties of the $\psi_{k}^{(-n_{1}-1,-n_{2}-1)}(x;N,M;\lambda_{1},\lambda_{2})$, we deduce that the $Q_{k}^{(n_{1},n_{2})}(z;N,M;\lambda_{1},\lambda_{2})$, $k=-n_{1}-1,-n_{2}-1,0,1,2,\ldots$, are orthogonal polynomials on ]1,1[]-1,1[ with respect to the measure
@@ C Explicit examples
Consider for instance the case where $N=M=3$. The seed function indices $\left(n_{1},n_{2}\right)$ of the two-step chain can be chosen in the set ($3\leq n_{2}<n_{1}<6$), so that $(n_{1},n_{2})\in\left\{\left(4,3\right),(5,4),(5,3)\right\}$. As for the possible values of the regularity parameters $\lambda_{n_{i}}^{\left(N,M\right)},\ i=1,2$, we get from Eq. (3.22)
$\lambda_{3}^{\left(3,3\right)}=3,\ \lambda_{4}^{\left(3,3\right)}=2,\ \lambda_{5}^{\left(3,3\right)}=1/6.\$(5.44)
@@
Figure 1: Potential $V^{(-5,-4)}(z;3,3;-1,1)$ in terms of $z$ for 1<z<1-1<z<1.
Figure 2: Potential $V^{(-5,-4)}(z;3,3;t,1)$ in terms of $z$ and $t$ for 1<z<1-1<z<1 and 2<t<0-2<t<0.
Figure 3: Potential $V^{(-5,-4)}(z;3,3;-1,\mu)$ in terms of $z$ and $\mu$ for 1<z<1-1<z<1 and $0<\mu<3$.
[…]
@@ A Regularity conditions
On assuming that the extension obtained at the (m1)(m-1)th step is regular, which imposes that $\max\left(N,M\right)\leq n_{m-1}<\cdots<n_{2}<n_{1}<N+M$ and definite conditions on $\lambda_{1},\lambda_{2},\ldots,\lambda_{m-1}$, we want to impose that the same is true at the $m$th step. For this to occur, $\psi^{(N_{m-1})}_{-n_{m}-1}$ must be in the disconjugacy sector of $V^{(N_{m-1})}$, i.e., $\max\left(N,M\right)\leq n_{m}<n_{m-1}<\cdots<n_{1}<N+M$, and, in addition, it must not change sign on $0,\pi/2[$, which leads to the condition
@@
$\qquad\times W\left(p_{n_{1}}^{(-N,-M)}(z;\lambda_{1}),p_{n_{2}}^{(-N,-M)}(z;\lambda_{2}),\ldots,p_{n_{m}}^{(-N,-M)}(z;\lambda_{m})\mid z\right),$ (6.4)
@@
$\qquad\times\frac{b_{n_{1}}^{\left(N,M\right)}...b_{n_{m}}^{\left(N,M\right)}}{2^{m\left(m-1\right)/2}\left(M-1\right)^{m-1}\left(M-2\right)^{m-2}...(M-m+1)}D(n_{1},...,n_{m}),$ (6.33)
@@
As we have assumed that the (m1m-1)-step chain $N_{m-1}=\left(-n_{1}-1,\ldots,-n_{{m-1}}-1\right)$ is regular at each level, the condition above then implies that
@@ B Rationally-extended potential
which turns out to be a rational extension of V(x;Nm,Mm)V(x;N-m,M-m),
@@
For $V^{\left(N_{m}\right)}\left(x;N,M;\lambda_{1},\ldots,\lambda_{m}\right)$ to be a confining potential on $0,\pi/2[$, we need to impose that $V\left(x;N-m,M-m\right)$ has this property, which is achieved for $N,M>m+1/2$, thence $N,M\geq m+1$. For such $N,M$ values, we note that $V^{\left(N_{m}\right)}\left(x;N,M;\lambda_{1},\ldots,\lambda_{m}\right)$ is actually strongly repulsive in both 0 and $\pi/2$, since the singularities are there of the type $g/x^{2}$ ($g\geq 3/4$) and $g/\left(\pi/2-x\right)^{2}$ ($g\geq 3/4$), respectively. This means that at each extremity, only one basis solution is quadratically integrable [33].
@@
$\qquad\times\frac{W(p_{n_{1}}^{(-N,-M)}(z;\lambda_{1}),\ldots,p_{n_{m}}^{(-N,-M)}(z;\lambda_{m}),(1-z)^{N}(1+z)^{M}P_{k}^{(N,M)}(z)\mid z)}{W\left(p_{n_{1}}^{(-N,-M)}(z;\lambda_{1}),p_{n_{2}}^{(-N,-M)}(z;\lambda_{2}),\ldots,p_{n_{m}}^{(-N,-M)}(z;\lambda_{m})\mid z\right)}$
@@
$\qquad\times\frac{\psi_{-1}(x;N-m,M-m)}{\det\left[\mathbf{R}\left(z;\lambda_{1},...,\lambda_{m}\right)\right]},$ (6.49)
@@
$\times\frac{W(p_{n_{1}}^{(-N,-M)}(z;\lambda_{1}),\ldots,\check{p}_{n_{i}}^{(-N,-M)}(z;\lambda_{i}),\ldots,p_{n_{m}}^{(-N,-M)}(z;\lambda_{m})\mid z)}{W\left(p_{n_{1}}^{(-N,-M)}(z;\lambda_{1}),\ldots,p_{n_{m}}^{(-N,-M)}(z;\lambda_{m})\mid z\right)},$ (6.53)
@@
where $Q^{(n_{1},n_{2};\ldots,n_{m})}_{-n_{i}-1}(x;N,M;\lambda_{1},\ldots,\lambda_{m})$ is obtained from the determinant of $\mathbf{R}(z;\lambda_{1},\ldots,\lambda_{m})$ by suppressing the $m^{\rm th}$ row and the $i^{\rm th}$ column, i.e., by a minor that we denote by $\det[\mathbf{R}(z;\lambda_{1},\ldots,\lambda_{m})]^{m,i}$,
@@
Due to the orthogonality properties of the $\psi_{k}^{(N_{m})}$, the polynomials $Q_{k}^{(n_{1},n_{2},\ldots,n_{m})}(z;N,M;\lambda_{1},\ldots,\lambda_{m})$, $k=-n_{1}-1$, …, $n_{m}-1,0,1,2,\ldots$, form a set of orthogonal polynomials on ]1,1[]-1,1[ with respect to the measure